OpenStudy (anonymous):

graph the function f(x)=6^x

OpenStudy (anonymous):

f(0) = 6^0 = 1 f(1) = 6^1 = 6 f(2) = 6^2 =36 http://www.wolframalpha.com/input/?i=f%28x%29%3D6^x